# Direct Preference Optimization Using Sparse Feature-Level Constraints

### v1

@@ Abstract
The alignment of large language models (LLMs) with human preferences remains a key challenge. While post-training techniques like Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO) have achieved notable success, they often experience computational inefficiencies and training instability. In this paper, we propose Feature-level constrained Preference Optimization (FPO), a novel method designed to simplify the alignment process while ensuring stability. FPO leverages pre-trained Sparse Autoencoders (SAEs) and introduces feature-level constraints, allowing for efficient, sparsity-enforced alignment. Our approach enjoys efficiency by using sparse features activated in a well-trained sparse autoencoder and the quality of sequential KL divergence by using the feature-level offline reference. Experimental results on benchmark datasets demonstrate that FPO achieves an above 5% absolute improvement in win rate with much lower computational cost compared to state-of-the-art baselines, making it a promising solution for efficient and controllable LLM alignments.
@@ 2 Preliminary
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right].$ (2)
@@
SimPO simplifies preference optimization by removing the need for a reference model and aligning rewards directly with the length-normalized log-likelihood of the policy model’s output. The SimPO objective can be formulated as:
$\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right],$ (3)
where $\gamma$ is a positive margin ensuring the reward for the preferred response exceeds that of the less preferred one by at least $\gamma$. SimPO’s key innovations are (1) eliminating the reference model and (2) incorporating a target reward margin $\gamma$. However, while SimPO is computationally efficient, the lack of reference control (Roy et al., 2021) results in instability. As shown by Liu et al. (2024), the reference model plays a crucial role in stabilizing training and improving performance.
@@
Token-Level Direct Preference Optimization (TDPO) refines the DPO framework by operating at the token level, accounting for the sequential nature of text generation. The TDPO objective function is defined as:
$\max_{\pi_{\theta}}\mathbb{E}_{x,y^{<t}\sim\mathcal{D},z\sim\pi_{\theta}(\cdot|[x,y^{<t}])}\left[A_{\pi_{\text{ref}}}([x,y^{<t}],z)-\beta D_{\text{KL}}(\pi_{\theta}(\cdot|[x,y^{<t}])\|\pi_{\text{ref}}(\cdot|[x,y^{<t}]))\right],$
where $A_{\pi_{\text{ref}}}([x,y^{<t}],z)$ is the token-level advantage function, and $D_{\text{KL}}(\pi_{1}||\pi_{2})$ denotes the KL divergence between $\pi_{1}$ and $\pi_{2}$. The first version of the loss function is given by:
$\mathcal{L}_{\text{TDPO}_{1}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\delta_{\text{TDPO}_{1}}(x,y_{w},y_{l})\right)\right],$ (4)
where $\delta_{\text{TDPO}_{1}}(x,y_{w},y_{l})$ is the difference in forward KL divergence between the preferred and less preferred completions:
$\delta_{\text{TDPO}_{1}}(x,y_{w},y_{l})=\beta D_{\text{TDPO}_{1}}\left(x,y_{l};\pi_{\text{ref}}\|\pi_{\theta}\right)-\beta D_{\text{TDPO}_{1}}\left(x,y_{w};\pi_{\text{ref}}\|\pi_{\theta}\right),$ (5)
and the sequential KL divergence between policy and reference output with sequence length $T$ is defined as $D_{\text{TDPO}}({x},{y};\pi_{\mathrm{ref}}\|\pi_{\theta})=\sum\limits_{t=1}^{T}D_{\mathrm{KL}}(\pi_{\mathrm{ref}}(\cdot|[{x},y^{<t}])\|\pi_{\theta}(\cdot|[{x},y^{<t}]))$. To further stabilize the gradient within the optimization, an improved loss function $\mathcal{L}_{\text{TDPO}_{2}}$ is given by replacing the regularization $\delta_{\text{TDPO}_{1}}$ with:
$\delta_{\text{TDPO}_{2}}(x,y_{w},y_{l})=\alpha\left(\beta D_{\text{TDPO}}\left(x,y_{l};\pi_{\text{ref}}\|\pi_{\theta}\right)-\mathrm{sg}\left(\beta D_{\text{TDPO}}\left(x,y_{w};\pi_{\text{ref}}\|\pi_{\theta}\right)\right)\right),$ (6)
@@ 3 Feature-Level Direct Preference Optimization
Method Reference Efficiency Constraint
SFT Free High Weak
DPO Offline High Weak
SimPO Free High Weak
TDPO Needed Low Strong / Dense
FPO(Ours) Offline High Strong / Sparse
@@
  • Interpretability methods such as SAE are widely used for interpreting the inner representations of LLMs due to their sparse and monosemantic activations Chen et al. (2017); Huben et al. (2024). However, this feature has not yet been applied in areas outside of interpretability.
DPO with Reference-base Target Margin.
To begin, we examine the loss functions of DPO and its enhanced variants, specifically SimPO and TDPO. By comparing Equation 2 and Equation 4, we notice that TDPO and DPO share an identical implicit reward difference term: $\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$. Essentially, TDPO can be viewed as an extension of DPO, where a KL constraint $\delta(x,y_{w},y_{l})$ is incorporated into the sigmoid function $\sigma(\cdot)$ in addition to the implicit reward difference. Taking a step further, we can isolate $\pi_{\text{ref}}$ from each implicit reward term:
$\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$ $=\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)$ (9)
$\quad-\underbrace{\beta\left(\log\pi_{\text{ref}}(y_{w}|x)-\log\pi_{\text{ref}}(y_{l}|x)\right)}_{\mathrel{\mathop{:}}=\gamma_{\text{ref}}}.$
We can see that Equation 9 shares a similar form with the reward difference calculation of SimPO in Equation 3. This similarity reveals that the reward difference in DPO can be interpreted as a combination of log probability difference with an adaptive margin $\gamma_{\text{ref}}$ from the reference model, whereas SimPO calculates the average log probability difference with a fixed margin. Based on the above observation, we can reframe the loss function of DPO and its two variants into a unified form:
$\mathcal{L}_{\text{PO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\underbrace{u(x,y_{w},y_{l})}_{\text{Log Probability Difference}}-\underbrace{\gamma(x,y_{w},y_{l})}_{\text{Margin}}-\underbrace{\delta(x,y_{w},y_{l})}_{\text{Constraint}}\right)\right].$ (10)
@@
TDPOi $\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)$ $\gamma_{\text{ref}}$ $\delta_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\text{TDPO}_{i}}}(x,y_{w},y_{l})$ KL Divergence
@@
$\frac{\beta}{|y_{w}|}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\frac{\beta}{|y_{l}|}\beta$ $\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}=\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)$ (11)
$\quad-\underbrace{\left(\frac{\beta}{|y_{w}|}\log\pi_{\text{ref}}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\text{ref}}(y_{l}|x)\right)}_{\mathrel{\mathop{:}}=\gamma_{\text{ref-LN}}}.$
Equation 11 suggests using average log probability difference as the Log Probability Difference (LPD) term and introducing an adaptive margin with length normalization as the Margin. The length-normalized margin $\gamma_{\text{ref-LN}}$ enhances the stability by using a reference model to calculate an adaptive margin for each preference pair. We consider an offline caching technique to minimize the computational overhead introduced by the reference model.
@@
Currently, the use of constraints $\delta(x,y_{w},y_{l})$ in alignment processes typically follows KL divergence-based approach shown in Equation 5 and 6. However, this method has a significant issue: for most LLMs, which generally have a very large vocabulary, where we assume the vocabulary size is $V$. For each batch with an input length of $T$, the resulting output probabilities have a size of $V\times T$. This work adopts Gemma (Lieberum et al., 2024), an advanced open-sourced LLM series, which has a massive vocabulary size of 265K. For an input length of 1024, this results in a probabilities matrix containing approximately 262M elements, which is nearly 1/10 the size of its 2B version model. Therefore, computing the KL divergence incurs considerable computational overhead to DPO-enhancing methods such as TDPO.
LLMs generate these sizable output probabilities by projecting their internal representations onto vocabulary space. In contrast to this, SAE is found to be capable of projecting these representations onto a sparse and more interpretable feature space. Motivated by the efficient nature of sparsity, we leverage the sparse feature activations from SAE to approximate the function that token-level probabilities serve in KL divergence. Specifically, for the output representation $h^{(t,\ell)}$ from layer $\ell$ of the model at position $t$, we can obtain its sparse activation $c^{(t,\ell)}$ using an SAE as described in Equation 7. Since KL divergence measures the difference between two probability distributions, we employ MSE as the loss to measure the discrepancy between the sparse activation from the two models. To further improve efficiency, instead of calculating the sum of token-wise discrepancy like TDPO, we first perform average pooling for the sparse activation across tokens and then calculate the MSE between pooled sparse activations, which gives us a more efficient sequential discrepancy:
@@
We have justified the implementation of the key components in Equation 9, which is a SimPO-like reward difference with a reference-based adaptive margin and a feature-level constraint. At first glance, the reference model appears to be deeply involved in both the calculation of the margin and the constraint, making its complete elimination challenging. Therefore, instead of directly removing the reference model, we propose a more appropriate approach: separating the computation of the reference model from the training process by computing its output offline. Offline computation means pre-calculating and caching the results related to the reference model needed for training and then reading them during the training loop. This approach allows us to free up the reference model during alignment with only a small and acceptable I/O demand.
To explore an implementation for Equation 10 that enjoys the advantages of SimPO, such as length normalization, while ensuring stability, first, we pre-compute and store the margin $\gamma_{\text{ref-LN}}$ using the length normalization for each preference pair. Since it is scalar, it only occupies $O(N)$ space to store it, where $N$ is the number of preference pairs. Next, for the feature-level constraint, we pre-compute and store the sparse activation of each sample in the training dataset following the computation in Equation 12. Consequently, we only need to pre-compute and store one sparse activation $\bar{c}^{\ell}_{\text{ref}}$ for each sample, which requires $O(2\cdot N\cdot k)$ space. This results in a significantly smaller space requirement compared to constraints used in TDPO, where the vocabulary size is $V$, for each batch with an input length of $T$, requiring a much larger space of $O(2\cdot N\cdot T\cdot V)$. By combining all the above results, we arrive at the loss function for FPO:
$\begin{split}\mathcal{L}_{\text{FPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\Big{[}\log\sigma\Big{(}&\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)\\ &-\gamma_{\text{ref-LN}}-\delta^{\ell}_{\text{FPO}}(x,y_{w},y_{l})\Big{)}\Big{]}\end{split}.$ (13)

### v4

# Constrain Alignment with Sparse Autoencoders

@@ Abstract
The alignment of large language models (LLMs) with human preferences remains a key challenge. While post-training techniques like Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO) have achieved notable success, they often experience computational inefficiencies and training instability. In this paper, we propose Feature-level constrained Preference Optimization (FPO), a novel method designed to simplify the alignment process while ensuring stability. FPO leverages pre-trained Sparse Autoencoders (SAEs) and introduces feature-level constraints, allowing for efficient, sparsity-enforced alignment. Our approach enjoys efficiency by using sparse features activated in a well-trained sparse autoencoder and the quality of sequential KL divergence by using the feature-level offline reference. Experimental results on benchmark datasets demonstrate that FPO achieves an above 5% absolute improvement in win rate with much lower computational cost compared to state-of-the-art baselines, making it a promising solution for efficient and controllable LLM alignments. Code is available at FeatureAlignment.
@@ 1 Introduction
Method Reference Efficiency Constraint
SFT Free High Weak
DPO Offline High Weak
SimPO Free High Weak
TDPO Needed Low Strong / Dense
FPO(Ours) Offline High Strong / Sparse
@@ 2 Preliminary
$\begin{split}\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(u(x,y_{w},y_{l})\right)\right],\\ u(x,y_{w},y_{l})=\beta(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}).\end{split}$
@@
SimPO simplifies DPO by removing the need for a reference model and aligning rewards directly with the length-normalized log-likelihood of the policy model’s output. The SimPO loss function can be formulated as:
$\begin{split}\mathcal{L}_{\text{SimPO}}(\pi_{\theta})&=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(u(x,y_{w},y_{l})\right)\right],\\ u(x,y_{w},y_{l})&=\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma.\end{split}$
where $\gamma$ is a positive margin ensuring the reward for the preferred response exceeds that of the less preferred one by at least $\gamma$. However, while SimPO is computationally efficient, the lack of reference control (Roy et al., 2021) results in instability, as the reference model can stabilize training and improving performance (Liu et al., 2024).
@@
Token-Level Direct Preference Optimization (TDPO) refines the DPO framework by operating at the token level, accounting for the sequential nature of text generation. The first version of TDPO loss function is given by:
$\small\begin{split}\mathcal{L}_{\text{TDPO}_{1}}(\pi_{\theta};\pi_{\text{ref}})&=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(u(x,y_{w},y_{l})\right)\right],\\ u(x,y_{w},y_{l})&=\\ \beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}&-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\delta_{\text{TDPO}_{1}}(x,y_{w},y_{l})\end{split}$ (2)
where $\delta_{\text{TDPO}_{1}}(x,y_{w},y_{l})$ is the KL divergence difference between the preferred and less preferred completions:
$\small\begin{split}\delta_{\text{TDPO}_{1}}(x,&y_{w},y_{l})=\\ \beta(D_{\text{TDPO}_{1}}&\left(x,y_{l};\pi_{\text{ref}}\|\pi_{\theta}\right)-D_{\text{TDPO}_{1}}\left(x,y_{w};\pi_{\text{ref}}\|\pi_{\theta}\right),\end{split}$ (3)
and the sequential KL divergence between policy and reference output with sequence length $T$ is defined as
$\small\begin{split}D_{\text{TDPO}}({x},{y};\pi_{\mathrm{ref}}\|\pi_{\theta})=\sum\limits_{t=1}^{T}D_{\mathrm{KL}}&(\pi_{\mathrm{ref}}(\cdot|[{x},y^{<t}])\|\pi_{\theta}(\cdot|[{x},y^{<t}]))\end{split}$
To further stabilize the gradient within the optimization, an improved loss function $\mathcal{L}_{\text{TDPO}_{2}}$ is given by replacing the regularization $\delta_{\text{TDPO}_{1}}$ with:
$\small\begin{split}\delta_{\text{TDPO}_{2}}\left(x,y_{w},y_{l}\right)=\alpha&\left(\beta D_{\text{TDPO}}\left(x,y_{l};\pi_{\text{ref}}\|\pi_{\theta}\right)\right.\\ &\left.-\mathrm{sg}\left(\beta D_{\text{TDPO}}\left(x,y_{w};\pi_{\text{ref}}\|\pi_{\theta}\right)\right)\right),\end{split}$ (4)
@@ 3 Feature-Level Direct Preference Optimization
  • Interpretability methods such as SAE are widely used for interpreting the inner representations of LLMs due to their sparse and monosemantic activations (Chen et al., 2017; Huben et al., 2024). However, this feature has not yet been applied in areas outside of interpretability.
@@
TDPOi $\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)$ $\gamma_{\text{ref}}$ $\delta_{{\text{TDPO}_{i}}}(x,y_{w},y_{l})$ KL Divergence
@@
DPO with Reference-base Target Margin.
To begin, we examine the loss functions of DPO and its enhanced variants, specifically SimPO and TDPO. By comparing Section 2 and Equation 2, we notice that TDPO and DPO share an identical implicit reward difference term: $\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$. Essentially, TDPO can be viewed as an extension of DPO, where a KL constraint $\delta(x,y_{w},y_{l})$ is incorporated into the sigmoid function $\sigma(\cdot)$ in addition to the implicit reward difference. Taking a step further, we can isolate $\pi_{\text{ref}}$ from each implicit reward term:
$\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$ $=$ (7)
$\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)$
$\quad-\underbrace{\beta\left(\log\pi_{\text{ref}}(y_{w}|x)-\log\pi_{\text{ref}}(y_{l}|x)\right)}_{\coloneq\gamma_{\text{ref}}}.$
We can see that Equation 7 shares a similar form with the reward difference calculation of SimPO in Section 2. This similarity reveals that the reward difference in DPO can be interpreted as a combination of log probability difference with an adaptive margin $\gamma_{\text{ref}}$ from the reference model, whereas SimPO calculates the average log probability difference with a fixed margin. Based on the above observation, we can reframe the loss function of DPO and its two variants into a unified form:
$\small\begin{split}\mathcal{L}_{\text{FPO}}(\pi_{\theta};\pi_{\text{ref}})&=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\Big{[}\log\sigma\Big{(}u(x,y_{w},y_{l})\Big{)}\Big{]},\\ u(x,y_{w},y_{l})&=\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)\\ &-\gamma_{\text{ref-LN}}-\delta^{\ell}_{\text{FPO}}(x,y_{w},y_{l}).\end{split}$ (8)
@@
$\frac{\beta}{|y_{w}|}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\frac{\beta}{|y_{l}|}\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$ (9)
$\quad-\left(\frac{\beta}{|y_{w}|}\log\pi_{\text{ref}}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\text{ref}}(y_{l}|x)\right)$
$=\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma_{\text{ref-LN}}.$
LABEL:eq:dpo_ln suggests using average log probability difference as the Log Probability Difference (LPD) term and introducing an adaptive margin with length normalization as the Margin. The length-normalized margin $\gamma_{\text{ref-LN}}$ enhances the stability by using a reference model to calculate an adaptive margin for each preference pair. We consider an offline caching technique to minimize the computational overhead introduced by the reference model.
@@
Currently, the use of constraints $\delta(x,y_{w},y_{l})$ in alignment processes typically follows KL divergence-based approach shown in Equation 3 and 4. However, this method has a significant issue: for most LLMs, which generally have a very large vocabulary, where we assume the vocabulary size is $V$. For each batch with an input length of $T$, the resulting output probabilities have a size of $V\times T$. This work adopts Gemma (Lieberum et al., 2024), an advanced open-sourced LLM series, which has a massive vocabulary size of 265K. For an input length of 1024, this results in a probabilities matrix containing approximately 262M elements, which is nearly 1/10 the size of its 2B version model. Therefore, computing the KL divergence incurs considerable computational overhead to DPO-enhancing methods such as TDPO.
LLMs generate these sizable output probabilities by projecting their internal representations onto vocabulary space. In contrast to this, SAE is found to be capable of projecting these representations onto a sparse feature space. Motivated by the efficient nature of sparsity, we leverage the sparse feature activations from SAE to approximate the function of KL divergence. Specifically, for the output representation $h^{(t,\ell)}$ from layer $\ell$ of the model at position $t$, we can obtain its sparse activation $c^{(t,\ell)}$ using an SAE as described in Equation 5. Since KL divergence measures the difference between two probability distributions, we employ MSE as the loss to measure the discrepancy between the sparse activation from the two models. To further improve efficiency, instead of calculating the sum of token-wise discrepancy like TDPO, we first perform average pooling for the sparse activation across tokens and then calculate the MSE between pooled sparse activations, which gives us a more efficient sequential discrepancy:
@@
We have justified the implementation of the key components in Equation 7, which is a SimPO-like reward difference with a reference-based adaptive margin and a feature-level constraint. At first glance, the reference model appears to be deeply involved in both the calculation of the margin and the constraint, making its complete elimination challenging.
Therefore, instead of directly removing the reference model, we propose a more appropriate approach: separating the computation of the reference model from the training process by computing its output offline. Offline computation means pre-calculating and caching the results related to the reference model needed for training and then reading them during the training loop. This approach allows us to free up the reference model during alignment with only a small and acceptable I/O demand.
To explore an implementation for Equation 8 that enjoys the advantages of SimPO, such as length normalization, while ensuring stability, first, we pre-compute and store the margin $\gamma_{\text{ref-LN}}$ using the length normalization for each preference pair. Since it is scalar, it only occupies $O({N})$ space to store it, where ${N}$ is the number of preference pairs. Next, for the feature-level constraint, we pre-compute and store the sparse activation of each sample in the training dataset following the computation in Equation 10. Consequently, we only need to pre-compute and store one sparse activation $\bar{c}^{\ell}_{\text{ref}}$ for each sample, which requires $O(2\cdot{N}\cdot k)$ space. This results in a significantly smaller space requirement compared to constraints used in TDPO, where the vocabulary size is $V$, for each batch with ${N}$ preference pairs, requiring a much larger space of $O(2{N}\cdot V)$. By combining all the above results, we arrive at the loss function for FPO:
$\small\begin{split}\mathcal{L}_{\text{FPO}}(\pi_{\theta};\pi_{\text{ref}})&=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\Big{[}\log\sigma\Big{(}u(x,y_{w},y_{l})\Big{)}\Big{]},\\ u(x,y_{w},y_{l})&=\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)\\ &-\gamma_{\text{ref-LN}}-\delta^{\ell}_{\text{FPO}}(x,y_{w},y_{l}).\end{split}$ (11)
Table 2: Left: Performance comparison of different methods for Gemma-2-2B and Gemma-2-9B across various benchmarks (AlpacaEval-2, Arena-Hard, and MT-Bench), compared to Supervised Fine-Tuning (SFT), DPO and variants. Length controlled Winning Rate: WR-L; Winning Rate: WR. Right: Comparison of FPO and other baseline methods in terms of the trade-off between Alignment Acc(accuracy) and Diversity $H$ i.e., Diversity (Entropy) on the UltraFeedback dataset.
Gemma-2-2B Gemma-2-9B
Method AlpacaEval-2 Arena-Hard MT-Bench AlpacaEval-2 Arena-Hard MT-Bench
FPO v.s. WR-L(%) WR (%) WR (%) $\Delta$ Score WR-L (%) WR (%) WR (%) $\Delta$ Score
SFT 54.7 55.1 53.2 +0.5 51.2 52.4 53.4 +0.3
DPO 51.7 50.8 51.6 +0.1 51.0 51.0 51.2 +0.1
TDPO-1 51.5 54.4 51.4 +0.3 50.8 50.2 51.8 +0.1
TDPO-2 50.9 54.0 50.6 +0.2 50.2 49.9 49.5 0.0
SimPO 51.1 52.2 51.4 +0.4 50.2 51.8 51.0 +0.2
Method Acc(%) $\uparrow$ $H$uparrow
DPO 59.9 1.66
TDPO-1 63.2 1.65
TDPO-2 64.2 1.68
SimPO 63.4 1.64
FPO 64.1 1.68